OpenStudy (anonymous):

I guess I can fan and Medal. But don't ask for it i'll give it to you after you've helped. Name the property illustrated by the equation. 5a + (-5a) = 0

OpenStudy (anonymous):

it's multiplication

OpenStudy (anonymous):

Apparently I forgot the answers too, god bless. @alibaby Gracias, So it's infact, D. Multiplicative Inverse?

OpenStudy (anonymous):

welcome

zepdrix (zepdrix):

Hey there :) Here are some useful Addition Properties:\[\large\rm a+(b+c)=(a+b)+c\qquad\qquad\text{Associativity}\]\[\large\rm a+b=b+a\qquad\qquad\quad~~\qquad\qquad Commutativity\]\[\large\rm x+(-x)=0\qquad\qquad\qquad\qquad\qquad Additive~Inverse\] Does your problem look like any of those? :)

zepdrix (zepdrix):

btw, we don't know what you mean by D, you didn't provide the options.

OpenStudy (anonymous):

Options: a. Additive Inverse c. Associative Property b. Additive Identity d. Multiplicative Inverse

OpenStudy (anonymous):

yes it's multiplication inverse

zepdrix (zepdrix):

Mm k :) Well I gave you some tips to look at, hopefully it helps. (Btw ali's suggestion is wrong)

OpenStudy (anonymous):

Examples: 6 miles a day for d days → 6d 15 minutes fewer than m minutes → m – 15 When you evaluate an algebraic expression, you find the value of the expression for a given value of the variable. To evaluate an expression, you substitute the value of the variable into the expression and simplify. It is important to follow the order of operations if the expression contains more than one operation. Example: Evaluate t + 6 when t = 22. t + 6 = 22 + 6 = 28 An algebraic expression contains terms. A variable term consists of a variable and a coefficient, like 3z. An algebraic expression always contains at least one variable term. An expression may also contain a constant term, which is a fixed value with no variable, like 5. You can use the properties of operations to simplify expressions with like terms. Example: Simplify 8t + 3 – 6 + 2t. 8t + 3 – 6 + 2t = (8 + 2)t + (3 – 6) = 10t – 3 You can apply the Distributive Property to expand expressions to simplify the expression. Example: 2(9n – 4) + 3 = (2)(9)n – (2)(4) + 3 = (18n – 8) + 3 = 18n + (–8) + 3 = 18n ­– 5 You can also use the Distributive Property to factor expressions